The survival rate during a risky operation for patients with no other hope of survival is 81%. What is the probability that exactly four of the next five patients survive this operation? (Give your answer correct to three decimal places.)
Answer #1

Solution

Given that ,

p =81%= 0.81

q = 1 - p = 1 - 0.81 = 0.19

n = 5

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X = 4) = (5 C 4) * 0.81 4 * (0.19)5- 4

probability =0.409